# Multi-Scale Protein Language Model for Unified Molecular Modeling

### v1

@@ Abstract
Protein language models have demonstrated significant potential in the field of protein engineering. However, current protein language models primarily operate at the residue scale, which limits their ability to provide information at the atom level. This limitation prevents us from fully exploiting the capabilities of protein language models for applications involving both proteins and small molecules. In this paper, we propose ms-ESM (multi-scale ESM), a novel approach that enables multi-scale unified molecular modeling. ms-ESM achieves this by pre-training on multi-scale code-switch protein sequences and utilizing a multi-scale position encoding to capture relationships among residues and atoms. Experimental results indicate that ms-ESM surpasses previous methods in protein-molecule tasks, demonstrating the full utilization of protein language models. Further investigations reveal that through unified molecular modeling, ms-ESM not only gains molecular knowledge but also retains its understanding of proteins.
@@ 1 Introduction
Figure 1: Overview of our multi-scale pre-training process. We mix protein datasets and molecular datasets to train ms-ESM. It is worth noting that the model’s input is either a molecule or a protein, but not paired protein-molecule data.
@@
To tackle the aforementioned challenges, in this paper, we propose ms-ESM (multi-scale ESM), which achieves multi-scale unified molecular modeling through (i) pre-training on multi-scale code-switch protein sequences and (ii) describing relationships among residues and atoms using a multi-scale position encoding.
First, drawing inspiration from the concept of multilingual code-switching in machine translation (Yang et al., 2020; Li et al., 2022a), [^2] They create sentences that switch between two or more languages to help the model learn multilingual knowledge. Yang et al. (2020) enhance multilingual model capabilities by substituting words in the source sentence with their translations in the target language. Similarly, Li et al. (2022a) improve these abilities by replacing a source word or phrase with its counterpart in a different language and then masking the corresponding target word. Collectively, these studies demonstrate that such code-switching techniques significantly strengthen the multilingual capabilities of machine translation models. ms-ESM introduces the learning of multi-scale knowledge through pre-training on multi-scale code-switch protein sequences, obtained by randomly unzipping protein residues into their constituent atoms. In such a scenario, ms-ESM can not only capture multi-scale aligned knowledge but also efficiently handle inputs at both the residue and atom scales simultaneously.
Second, ms-ESM employs a multi-scale position encoding to comprehensively differentiate between residues and atoms within the code-switch protein sequence. At the residue scale, we extend the original position encoding used in ESM to align with the current best practices in handling pure residue sequences, thereby avoiding ambiguous positional information across different scales, including atom-to-atom, residue-to-residue, and residue-to-atom relationships. At the atom scale, to describe the relationships among unzipped atoms, we employ a spatial distance matrix that directly encodes their 3D positions. With this approach, we can effectively describe all relationships among the entities within the code-switch sequence.
The improved experiment results demonstrate that ms-ESM surpasses previous methods in protein-molecule tasks, indicating the full utilization of protein language models. The solid performance in protein tasks suggests that ms-ESM, facilitated by the novel unified molecular modeling we first proposed, acquires molecular knowledge without sacrificing its understanding of proteins. Additionally, when applying ms-ESM to standard molecular benchmarks, it also outperforms several molecule-specific models. These findings clearly highlight the potential of unified molecular modeling.
@@
2 Proposed Method: ms-ESM
In this section, we will describe our multi-scale pre-training model, i.e., ms-ESM, in detail. Inspired by the concept of multi-lingual code-switching methods, ms-ESM initially generates multi-scale code-switch protein sequences by randomly unzipping partial residues. Through training on these sequences with carefully designed multi-scale position encoding, ms-ESM demonstrates its efficacy at both the residue and atom scales. When addressing protein-molecule tasks, i.e., tasks involving both proteins and small molecules, ms-ESM does not require any additional models and can fully leverage the potential of pre-training.
Specifically, in Section 2.1, we introduce the overall objective of training ms-ESM. Subsequently, in Section 2.2, we delve into the details of constructing a code-switch protein sequence and implementing the multi-scale pre-training approach. To describe the complicated position relationships within the code-switch sequence, we present our design of a multi-scale position encoding in Section 2.3.
@@ 2.1 Overview
We start with an overview of our multi-scale pre-training model, i.e., ms-ESM (Figure 1). Briefly, the total objective of our pre-training can be expressed as the following loss function:
$\mathcal{L}_{\theta}=\sum_{X_{i}\in D}$ $\mathcal{L}_{\textsc{mlm}}(\bar{X}_{i},E_{i};\theta)+\mathcal{L}_{\textsc{pdr}}(\bar{X}_{i},E_{i};\theta)$
$=\sum_{X_{i}\in D}$ $\mathcal{L}_{\textsc{mlm}}(\textsc{Unzip}(X_{i}),\textsc{mspe}(X_{i});\theta)~{}+$
@@
For each data $X_{i}$ in dataset $D$, we first create its code-switch sequence $\bar{X}_{i}$ by unzipping partial residues. Using the code-switch sequence, we employ Masked Language Modeling (MLM) and Pair-wise Distance Recovery (PDR) as pre-training tasks. We discuss the details of $\bar{X}_{i}$, $\mathcal{L}_{\textsc{MLM}}$, and $\mathcal{L}_{\textsc{PDR}}$ in Section 2.2. To account for the coexistence of residues and atoms in the sequence, we propose a Multi-Scale Position Encoding (MSPE) $E_{i}$ to describe the complicated position relationship within $\bar{X}_{i}$ (Section 2.3). We show more details of ms-ESM, including the parameterization of $\theta$ in Section 2.4. Notably, since we utilize molecule data in pre-training, ms-ESM can accept either proteins or molecules as inputs.
@@ 2.2 Multi-scale Pre-training
Specifically, in the residue scale, a protein $X$ can be seen as a sequence of $L$ residues, i.e., $X=(r_{1},\cdots,r_{i},\cdots,r_{L})$. Each residue $r_{i}$ further consists of a specific set of $N$ atoms $A_{i}=\{a_{i}^{1},\cdots,a_{i}^{N}\}$. To construct a code-switch protein sequence $\bar{X}$, we randomly select a group of residues and insert their corresponding atoms into $X$, which is essentially an unzipping process. For each unzipped residue, we provide the model with structural information of the residue at the atomic scale, thus offering the model very diverse structural knowledge. In particular, during the unzipping process, we assign a sequential order to the unzipped atoms. Here, we take the case of unzipping a single residue as an example, whereas in actual modeling, multiple residues can be unzipped. After inserting the atom set $A_{i}$ into $X$, i.e., unzipping the residue $r_{i}$, we obtain a code-switch sequence
@@
We also employ PDR as another pre-training task. Briefly, we use corrupted atoms as model input and ask model to recover the accurate Euclidean distances between these atoms. We corrupt the atoms by adding noises to their coordinates. Specifically, we replace the ground-truth coordinate with a randomly selected position that is within a certain range (Euclidean distances $<\epsilon$, Appendix A) of the true coordinate. Models are required to reconstruct the actual distances based on the corrupted coordinates. To avoid introducing residue-residue interactions that are very different from the interactions in small molecules, we only calculate PDR within residues, which can also make ms-ESM learn very diverse structural knowledge of residues.
@@
Notably, when training ms-ESM, we mix up a protein dataset $D_{p}$ and a molecule dataset $D_{m}$ as the final dataset, i.e., $D=D_{p}\cup D_{m}$. For a molecule from $D_{m}$, as it consists solely of atoms, its code-switch sequence $\bar{X}$ is the ordered set of all its atoms $\bar{A}$, and it does not have any residues, i.e., $\bar{R}=\emptyset$.
@@ 2.3 Multi-scale Position Encoding
In this section, we design a multi-scale position encoding $E$ to encode the positional relationships within a code-switch sequence. Specifically, $E$ contains a residue scale position encoding $E^{R}$ and an atom scale position encoding $E^{A}$, i.e., $E=(E^{R},E^{A})$. For $E^{R}$, we carefully extend an existing encoding method, allowing it to encode relationships from residues to atoms, while maintaining consistency with the original encoding when handling pure residue sequences. For $E^{A}$, to capture the relationships among atoms, we directly encode their 3D positions using a spatial distance matrix. The multi-scale encoding approach ensures that no ambiguous positional relationships affect the pre-training, enabling ms-ESM to perform effectively in both scales. Figure 3 illustrates the framework of our multi-scale position encoding. We will provide detailed explanations for each of them in the following paragraphs.
@@ 2.4 Integrating Multi-scale PE into Transformer
The parameterization $\theta$ of ms-ESM is slightly different from the original Transformer architecture proposed by Vaswani et al. (2017). To be specific, we begin by substituting the sinusoidal encoding in the Transformer with our residue scale position encoding $E^{R}$. For the atom scale position encoding $E^{A}$, we treat it as the bias term of self-attention layers (Luo et al., 2022; Zhou et al., 2023). The self-attention in ms-ESM can be calculated like:
@@
where $Q,K,V$ are the query, key, and value corresponding to $\bar{X}$. We refer readers to Vaswani et al. (2017) for more details of the original Transformer. With only slight modifications to the original Transformer architecture, ms-ESM is capable of simultaneously processing residues and atoms, making it a versatile model for various downstream tasks. Moreover, ms-ESM shows great compatibility with existing pre-training model, e.g., ESM series, which allows us to bulid up a better model based on previous study more easily.
Table 1: Performance comparison on Enzyme-Substrate Affinity Regression (ESAR) task and Enzyme-Substrate Pair Classification (ESPC) task. ms-ESM outperforms other models and achieves the state-of-the-art results, which indicates that ms-ESM operate at both the residue and atom scales successfully and our unified modeling harness the full potential of PLMs.
@@
$\text{Ours}_{\text{ XGBoost}}$ ms-ESM 35M ms-ESM 35M 0.623 0.548 0.742 90.6% 0.750 0.943
$\text{Ours}_{\text{ ProSmith}}$ ms-ESM 35M ms-ESM 35M 0.599 0.566 0.753 91.8% 0.781 0.954
Table 2: Performance comparison on drug-target affinity regression task. ms-ESM achieves the state-of-the-art results on most metrics.
@@
$\text{Ours}_{\text{ XGBoost}}$ ms-ESM 35M ms-ESM 35M 0.248 0.889 0.668
$\text{Ours}_{\text{ ProSmith}}$ ms-ESM 35M ms-ESM 35M 0.191 0.906 0.759
@@ 3 Experiments
To verify the effectiveness of the multi-scale unified pre-training, we evaluated the performance of ms-ESM in various protein-molecule tasks, i.e., tasks involving both proteins and small molecules (Section 3.1). The improved results indicate that our unified molecular modeling approach can take good advantage of the full potential of protein language models.
We also conduct ablation studies on position encoding, training data, and pre-training objectives to discuss how these elements affect the performance of our model (Section 3.2). Collectively, these studies demonstrate that each of these components plays a crucial role in ms-ESM.
In addition, we use multiple protein tasks to evaluate ms-ESM’s understanding of proteins (Section 3.3). The solid performance in these protein tasks suggests that ms-ESM does not sacrifice its protein understanding at both local and global level.
Further applying ms-ESM to standard molecular benchmarks (Section 3.4), we observe that ms-ESM can outperform several molecule-specific methods.
The detailed pre-training configuration, including pre-training datasets and hyperparameters, e.g., the unzip ratio of residues, can be found in Appendix A.
@@ 3.1 Main Results
We use tasks involving both proteins and molecules to prove that ms-ESM can operate at both residue and atom scales and our unified molecular modeling approach can exploit the full potential of PLMs.
@@
For protein-molecule tasks, we follow the benchmark protocol from ProSmith (Kroll et al., 2023b) to evaluate ms-ESM on three tasks, including enzyme-substrate affinity regression, drug-target affinity regression, and enzyme-substrate pair classification. Specifically, each task provides the protein residue sequence and the molecule SMILES string as input and requires models to determine whether the protein-molecule pair exhibits high affinity. Since our ms-ESM cannot directly process SMILES strings, we initially employ RDKit (Landrum et al., 2013) to generate the corresponding molecule conformations based on the SMILES representation. Subsequently, we extract the atom sequence and atom scale position encoding for ms-ESM. For additional fine-tuning details, please refer to Appendix B.1.
@@
We compare ms-ESM with multiple baselines on each tasks, including both supervised and pre-training baselines. For each baseline, we list their protein pre-training model and molecule pre-training model in corresponding tables. More details of each baseline can be seen in corresponding papers. Since ProSmith (Kroll et al., 2023b) is the only framework that offers a method to integrate protein pre-training models and molecule pre-training models, we adopt their framework and replace both the protein and molecule models with ms-ESM to ensure a fair comparison. We also provide an XGBoost (Chen & Guestrin, 2016) variant of ProSmith, which takes the concatenation of protein and molecule representation as features and can directly assess the compatibility of the two representations

### v2

@@ Abstract
Protein language models have demonstrated significant potential in the field of protein engineering. However, current protein language models primarily operate at the residue scale, which limits their ability to provide information at the atom level. This limitation prevents us from fully exploiting the capabilities of protein language models for applications involving both proteins and small molecules. In this paper, we propose ESM-AA (ESM All-Atom), a novel approach that enables atom-scale and residue-scale unified molecular modeling. ESM-AA achieves this by pre-training on multi-scale code-switch protein sequences and utilizing a multi-scale position encoding to capture relationships among residues and atoms. Experimental results indicate that ESM-AA surpasses previous methods in protein-molecule tasks, demonstrating the full utilization of protein language models. Further investigations reveal that through unified molecular modeling, ESM-AA not only gains molecular knowledge but also retains its understanding of proteins.
@@ 1 Introduction
Figure 1: Overview of our multi-scale pre-training process. We mix protein datasets and molecular datasets to train ESM-AA. It is worth noting that the model’s input is either a molecule or a protein, but not paired protein-molecule data.
@@
To tackle the aforementioned challenges, in this paper, we propose ESM-AA (ESM All-Atom), which achieves multi-scale unified molecular modeling through (i) pre-training on multi-scale code-switch protein sequences and (ii) describing relationships among residues and atoms using a multi-scale position encoding.
First, drawing inspiration from the concept of multilingual code-switching in machine translation (Yang et al., 2020; Li et al., 2022a), [^2] They create sentences that switch between two or more languages to help the model learn multilingual knowledge. Yang et al. (2020) enhance multilingual model capabilities by substituting words in the source sentence with their translations in the target language. Similarly, Li et al. (2022a) improve these abilities by replacing a source word or phrase with its counterpart in a different language and then masking the corresponding target word. Collectively, these studies demonstrate that such code-switching techniques significantly strengthen the multilingual capabilities of machine translation models. ESM-AA introduces the concept of learning multi-scale knowledge by pre-training on code-switch protein sequences. These sequences are a hybrid of sequence and structure data, derived from randomly unzipping protein residues into their constituent atoms and assigning coordinates to each unzipped atom. In such a scenario, ESM-AA can not only capture multi-scale aligned knowledge but also efficiently handle residue sequences and atomic coordinates.
Second, ESM-AA employs a multi-scale position encoding to comprehensively differentiate between residues and atoms within the code-switch protein sequence. At the residue scale, we extend the original position encoding used in ESM to align with the current best practices in handling pure residue sequences, thereby avoiding ambiguous positional information across different scales, including atom-to-atom, residue-to-residue, and residue-to-atom relationships. At the atom scale, to describe the relationships among unzipped atoms, we employ a spatial distance matrix that directly encodes their 3D positions. With this approach, we can effectively describe all relationships among the entities within the code-switch sequence.
We pre-train ESM-AA using a mixture of protein and small molecule data, and fine-tune it on a diverse set of benchmarks for evaluation. The improved experiment results demonstrate that ESM-AA surpasses previous methods in protein-molecule tasks, indicating the full utilization of protein language models. The solid performance in protein tasks suggests that ESM-AA, facilitated by the novel unified molecular modeling we first proposed, acquires molecular knowledge without sacrificing its understanding of proteins. Additionally, when applying ESM-AA to standard molecular benchmarks, it also outperforms several molecule-specific models. These findings clearly highlight the potential of unified molecular modeling.
@@
2 Proposed Method: ESM-AA
In this section, we will describe our multi-scale pre-training model, i.e., ESM-AA, in detail. Due to the vast number of atoms in a protein molecule, it is impossible to simultaneously input all atomic information of a protein into the model. Inspired by the concept of multi-lingual code-switching methods, ESM-AA initially generates multi-scale code-switch protein sequences by randomly unzipping partial residues. Through training on these sequences with carefully designed multi-scale position encoding, ESM-AA demonstrates its efficacy at both the residue and atom scales. When addressing protein-molecule tasks, i.e., tasks involving both proteins and small molecules, ESM-AA does not require any additional models and can fully leverage the potential of pre-training.
Specifically, in Section 2.1, we introduce the overall objective of training ESM-AA. Subsequently, in Section 2.2, we delve into the details of constructing a code-switch protein sequence and implementing the multi-scale pre-training approach. To describe the complicated position relationships within the code-switch sequence, we present our design of a multi-scale position encoding in Section 2.3.
@@ 2.1 Overview
We start with an overview of our multi-scale pre-training model, i.e., ESM-AA (Figure 1). Briefly, the total objective of our pre-training can be expressed as the following loss function:
$\mathcal{L}_{\theta}=\sum_{X_{i}\in B}$ $\mathcal{L}_{\textsc{mlm}}(\bar{X}_{i},E_{i};\theta)+\mathcal{L}_{\textsc{pdr}}(\bar{X}_{i},E_{i};\theta)$
$=\sum_{X_{i}\in B}$ $\mathcal{L}_{\textsc{mlm}}(\textsc{Unzip}(X_{i}),\textsc{mspe}(X_{i});\theta)~{}+$
@@
where $B$ is a batch of data sampled from the dataset $D$. For each data $X_{i}$ in dataset $D$, we first create its code-switch sequence $\bar{X}_{i}$ by unzipping partial residues. Using the code-switch sequence, we employ Masked Language Modeling (MLM) and Pair-wise Distance Recovery (PDR) as pre-training tasks. We discuss the details of $\bar{X}_{i}$, $\mathcal{L}_{\textsc{MLM}}$, and $\mathcal{L}_{\textsc{PDR}}$ in Section 2.2. To account for the coexistence of residues and atoms in the sequence, we propose a Multi-Scale Position Encoding (MSPE) $E_{i}$ to describe the complicated position relationship within $\bar{X}_{i}$ (Section 2.3). We show more details of ESM-AA, including the parameterization of $\theta$ in Section 2.4. Notably, since we utilize molecule data in pre-training, ESM-AA can accept either proteins or molecules as inputs.
@@ 2.2 Multi-scale Pre-training
Specifically, in the residue scale, a protein $X$ can be seen as a sequence of $L$ residues, i.e., $X=(r_{1},\cdots,r_{i},\cdots,r_{L})$. Each residue $r_{i}$ further consists of a specific set of $N$ atoms $A_{i}=\{a_{i}^{1},\cdots,a_{i}^{N}\}$. To construct a code-switch protein sequence $\bar{X}$, we randomly select a group of residues and insert their corresponding atoms into $X$, which is essentially an unzipping process. For each unzipped residue, we provide the model with structural information of the residue at the atomic scale, i.e., atomic coordinates, thus offering the model very diverse structural knowledge. In particular, during the unzipping process, we assign a sequential order to the unzipped atoms. Here, we take the case of unzipping a single residue as an example, whereas in actual modeling, multiple residues can be unzipped. After inserting the atom set $A_{i}$ into $X$, i.e., unzipping the residue $r_{i}$, we obtain a code-switch sequence
@@
We also employ PDR as another pre-training task. Briefly, we use corrupted atoms as model input and ask model to recover the accurate Euclidean distances between these atoms. We corrupt the atoms by adding noises to their coordinates. Specifically, we replace the ground-truth coordinate with a randomly selected position that is within a certain range (Euclidean distances $<\epsilon$, Appendix A) of the true coordinate. Models are required to reconstruct the actual distances based on the corrupted coordinates. To avoid introducing residue-residue interactions that are very different from the interactions in small molecules, we only calculate PDR within residues, which can also make ESM-AA learn very diverse structural knowledge of residues.
@@
Notably, when training ESM-AA, we mix up a protein dataset $D_{p}$ and a molecule dataset $D_{m}$ as the final dataset, i.e., $D=D_{p}\cup D_{m}$. For a molecule from $D_{m}$, as it consists solely of atoms, its code-switch sequence $\bar{X}$ is the ordered set of all its atoms $\bar{A}$, and it does not have any residues, i.e., $\bar{R}=\emptyset$.
@@ 2.3 Multi-scale Position Encoding
In this section, we design a multi-scale position encoding $E$ to encode the positional relationships within a code-switch sequence. Specifically, $E$ contains a residue scale position encoding $E^{R}$ and an atom scale position encoding $E^{A}$, i.e., $E=(E^{R},E^{A})$. For $E^{R}$, we carefully extend an existing encoding method, allowing it to encode relationships from residues to atoms, while maintaining consistency with the original encoding when handling pure residue sequences. For $E^{A}$, to capture the relationships among atoms, we directly encode their 3D positions using a spatial distance matrix. The multi-scale encoding approach ensures that no ambiguous positional relationships affect the pre-training, enabling ESM-AA to perform effectively in both scales. Figure 3 illustrates the framework of our multi-scale position encoding. We will provide detailed explanations for each of them in the following paragraphs.
@@ 2.4 Integrating Multi-scale PE into Transformer
The parameterization $\theta$ of ESM-AA is slightly different from the original Transformer architecture proposed by Vaswani et al. (2017). To be specific, we begin by substituting the sinusoidal encoding in the Transformer with our residue scale position encoding $E^{R}$. For the atom scale position encoding $E^{A}$, we treat it as the bias term of self-attention layers (Luo et al., 2022; Zhou et al., 2023). The self-attention in ESM-AA can be calculated like:
@@
where $Q,K,V$ are the query, key, and value corresponding to $\bar{X}$. We refer readers to Vaswani et al. (2017) for more details of the original Transformer. With only slight modifications to the original Transformer architecture, ESM-AA is capable of simultaneously processing residues and atoms, making it a versatile model for various downstream tasks. Moreover, ESM-AA shows great compatibility with existing pre-training model, e.g., ESM series, which allows us to bulid up a better model based on previous study more easily.
Table 1: Performance comparison on Enzyme-Substrate Affinity Regression (ESAR) task and Enzyme-Substrate Pair Classification (ESPC) task. ESM-AA outperforms other models and achieves the state-of-the-art results, which indicates that ESM-AA operate at both the residue and atom scales successfully and our unified modeling harness the full potential of PLMs.
@@
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.623 0.548 0.742 90.6% 0.750 0.943
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.599 0.566 0.753 91.8% 0.781 0.954
Table 2: Performance comparison on drug-target affinity regression task. ESM-AA achieves the state-of-the-art results on most metrics.
@@
$\text{Ours}_{\text{ XGBoost}}$ ESM-AA 35M ESM-AA 35M 0.248 0.889 0.668
$\text{Ours}_{\text{ ProSmith}}$ ESM-AA 35M ESM-AA 35M 0.191 0.906 0.759
@@ 3 Experiments
We pre-train ESM-AA on mixed data of proteins and small molecules. For the proteins, we construct code-switch sequences that contain both sequence and structural information, as described in Section 2.2. We fine-tune and evaluate ESM-AA across diverse benchmarks and verify the contribution of each component through ablation experiments. Finally, a visualization analysis is included to explain the advantages of unified modeling.
3.1 Pre-training Configuration
Datasets
We pre-train using a dataset that includes both protein and molecule data, specifically selecting those with structural details such as atom coordinates for encoding Euclidean distances and recovering pair-wise distances. For the protein dataset, we use AlphaFold DB (Varadi et al., 2022) dataset, which contains 8M protein sequences and structures predicted by AlphaFold2 (Jumper et al., 2021) with high confidence (pLDDT $>$ 90). For the molecule dataset, we use the dataset provided by Zhou et al. (2023), which contains 19M molecules and 209M conformations generated by ETKGD (Riniker & Landrum, 2015) and Merck Molecular Force Field (Halgren, 1996).
Hyperparameters
We implement ESM-AA using 12 stacked Transformer layers, each with 20 attention heads, as discussed in Section 2.4. The model dimension and feedforward dimension of each Transformer layer are 480 and 1920. We use Adam (Kingma & Ba, 2014) and polynomial learning rate scheduler to train ESM-AA and set the learning rate 4e-4, weight decay 1e-2, warmup step 5000. The total training step is 300K and each batch has 256K tokens at maximum. We train ESM-AA on 16 NVIDIA A100 GPU cards for 3 days. ESM-AA is compatible with ESM series, so we load a ESM-2 checkpoint as the initialization of ESM-AA. When pre-training, 1.0% of residues are unzipped as the pre-training setting, which makes the unzipped protein sequence 1.08 times longer than before on average. Thus we make an adjustment to the maximum sequence length permissible for ESM-AA, transitioning from ESM-2’s 1024 to 2048. Table 5 provides a complete list of hyperparameters.
@@ 3.2 Main Results
We use tasks involving both proteins and molecules to prove that ESM-AA can operate at both residue and atom scales and our unified molecular modeling approach can exploit the full potential of PLMs.
@@
For protein-molecule tasks, we follow the benchmark protocol from ProSmith (Kroll et al., 2023b) to evaluate ESM-AA on three tasks, including enzyme-substrate affinity regression, drug-target affinity regression, and enzyme-substrate pair classification. Specifically, each task provides the protein residue sequence and the molecule SMILES string as input and requires models to determine whether the protein-molecule pair exhibits high affinity. Since our ESM-AA cannot directly process SMILES strings, we initially employ RDKit (Landrum et al., 2013) to generate the corresponding molecule conformations based on the SMILES representation. Subsequently, we extract the atom sequence and atom scale position encoding for ESM-AA. For additional fine-tuning details (datasets and hyperparameters), please refer to Appendix B.1.
@@
We compare ESM-AA with multiple baselines on each tasks, including both supervised and pre-training baselines. For each baseline, we list their protein pre-training model and molecule pre-training model in corresponding tables. More details of each baseline can be seen in corresponding papers. We also use the standard framework provided by ProSmith for evaluating ESM-AA to ensure a fair comparison. Specifically, the framework contains three main modules, i.e., molecule encoder, protein encoder, and fusion block. Two encoders extract features from proteins and molecules severally. The fusion block is a Transformer model, which is responsible for fusing protein and molecule features. The fused features are further used to regress the affinity values or predict binary affinity. We compare performance by replacing encoders with different pre-trained models (ESM-AA, ESM-2, Uni-Mol). We also provide the results of an XGBoost (Chen & Guestrin, 2016) variant of ProSmith, which removes the fusion block and uses simple concatenation for feature fusing and can directly assess the compatibility of the two representations. Note that we freeze both encoders in the experiments as suggested by ProSmith. We turn off the unzip operation when performing fine-tuning.